find y as a function of x. The constant C is a positive number.
ln(y-1)+ln(y+1)=-x+c

Answer by sdmmadam@yahoo.com(530)

find y as a function of x. The constant C is a positive number.
ln(y-1)+ln(y+1)=-x+c
ln(y-1)+ln(y+1)=-x+c ----(1)
log[(y-1)(y+1)] = (-x+c)
(using formula loga +logb =log(ab)all to the same base 10)
log(y^2-1)=(-x+c)
(y^2-1)= 10^(-x+c)
Using definition: Given N>0,
log(N) to a given base b is the power p to which the base has to be raised to give the number. That is N =b^p
y^2 = [10^(c-x)+1]
y = (+ or minus)sqrt[10^(c-x)+1]